"3" x^2+16=x+4

"3"


1
Expert's answer
2022-04-14T00:55:00-0400

"3x^2+16=x+4\Rightarrow 3x^2-x+12=0\\D=1-4\cdot 3\cdot 12=-143<0\\No\,\,solutions"